You want to know the percentage of utility companies that earned revenue greater than 94 million dollars. If the mean revenue was 60 million dollars and the data has a standard deviation of 15 million, find the percentage. Assume that the distribution is normal. Round your answer to the nearest hundredth.
Answer #1

Solution :

Given that ,

mean = = 60

standard deviation = =15

P(x >94 ) = 1 - P(x<94 )

= 1 - P[(x -) / < (94 -60) / 15]

= 1 - P(z <2.27 )

Using z table

= 1 - 0.9884

= 0.0116

answer=1.16%